# Scaling and Predictability in Surface Quasi-Geostrophic Turbulence

### v1

@@ 1 Introduction
In this work, we consider the surface quasi-geostrophic (SQG) equation, a model that describes the flow governed by the conservation of buoyancy at the surface of a rotating, stratified fluid (Blumen, 1978; Pierrehumbert et al., 1994). Beyond its geophysical relevance, the SQG model has gained attention in the fluid dynamics community due to its striking similarities to three-dimensional (3D) Navier-Stokes (NS) turbulence while keeping some properties of two-dimensional (2D) flows. In particular, SQG has two inviscid quadratic invariants, similarly to 2D turbulence (Celani et al., 2004; Lapeyre, 2017; Valade et al., 2024) with one of the two, the surface potential energy, which displays, in the presence of forcing and dissipation, a direct cascade à la Kolmogorov towards the small scales similar to 3D turbulence (Valadão et al., 2024b). Despite this similarity, previous numerical studies (Pierrehumbert et al., 1994; Ohkitani & Yamada, 1997; Sukhatme & Pierrehumbert, 2002) reported that the scaling exponent of the spectrum of the surface potential energy deviates from the Kolmogorov value $-5/3$ predicted by dimensional arguments.
Based on very high-resolution direct numerical simulations of the SQG model, we study the statistical properties of the direct cascade. At large Reynolds number, we observe the recovery of the Kolmogorov scaling in the surface potential energy spectrum. Further, we measure the finite-time distribution of the Lyapunov exponent as a function of Re and we find an anomalous scaling law in which the Lyapunov exponent grows faster than what is predicted on dimensional grounds, similar to what was observed in 3D NS turbulence (Boffetta & Musacchio, 2017). Despite this anomaly, we find that the distribution of the finite-time Lyapunov exponents follow an almost universal function, independent of the Reynolds number of the flow.
@@ 2 The Surface Quasi-geostrophic model
The SQG equation describes the evolution of the surface buoyancy field $\theta({\bm{x}},t)$, governed by the following equation (Pierrehumbert et al., 1994):
$\partial_{t}\theta+{\bf v}\cdot{\bm{\nabla}}\theta=\nu\nabla^{2}\theta-\mu\nabla^{-2}\theta+f\;,$ (1)
where $f(\bm{x},t)$ represents a forcing, $\nu$ is a diffusion coefficient, and $\mu$ accounts for large-scale friction. We remark that most oceanic and atmospheric applications of SQG disregard dissipation terms as the model is designed to capture large-scale dynamics (Juckes, 1994; Lapeyre & Klein, 2006; Vallis, 2017; Siegelman et al., 2022). On the contrary, when studied numerically as a turbulent system, the model requires dissipative terms to avoid the small and large scale accumulation of the energy injected by the forcing term and to reach a statistically stationary state.
The incompressibility condition $\bm{\nabla}\cdot\bf v=0$ can be enforced through the definition of a stream function $\psi(\bm{x},t)$, such that ${\bf v}({\bm{x}},t)=(-\nabla_{y}\psi,\nabla_{x}\psi)$. Finally, the relation between the buoyancy field and the stream function is given by $\psi=\nabla^{-1}\theta$ or, in Fourier space, $\hat{\psi}=\hat{\theta}/k$ (where $k\equiv|\bm{k}|$) and consequently the velocity field can be written in terms of buoyancy as
@@
from which one observes that the buoyancy field is dimensionally a velocity.
@@
If one assumes that the forcing $f$ is active on a characteristic scale $\ell_{f}$, a turbulent state can develop with a double cascade phenomenology (Blumen, 1978; Pierrehumbert et al., 1994). Within this scenario, SPE is primarily transferred from the forcing scale to smaller ones ($\ell<\ell_{f}$), producing the direct cascade which is eventually dissipated by viscosity at the diffusive scale $\ell_{\nu}$. Meanwhile, VIE undergoes an inverse cascade, transferring energy to scales larger than the forcing ($\ell>\ell_{f}$) until it is dissipated at the friction scale $\ell_{\mu}$. In the statistically stationary state, the SPE and VIE balances are given by
@@
Under the assumptions of statistical homogeneity and isotropy, Blumen (Blumen, 1978) first predicted the power-law behavior of the energy spectrum $E(k)\equiv\langle|\hat{\theta}({\bm{k}})|^{2}\rangle/2$ for sufficiently large scale separations $\ell_{\nu}\ll\ell_{f}\ll\ell_{\mu}$. In this case, the spectral energy densities follow
@@
Note that in (10) Re is based on $\varepsilon_{I}$ and it is therefore defined a priori. Alternatively, we could use the VIE dissipation $\varepsilon_{\nu}$ resulting in a slightly smaller value of Re, as discussed below. We also remark that the scaling laws in (7-8) can also be obtained from the analogous of the exact four-fifths law of turbulence and by assuming self-similarity of the statistics (Frisch, 1995; Valadão et al., 2024b; Valade et al., 2025).
@@ 3 Numerical simulations and results
We explore the statistical properties and the Eulerian predictability of the direct cascade in SQG at different Reynolds numbers by numerically integrating (1) at high resolution with a pseudo-spectral, GPU accelerated code. Simulations are performed in a square domain of size $L_{x}=L_{y}=2\pi$ with periodic boundary conditions, using a regular grid with resolution $N\times N$ ranging from $N=1024$ to $N=16384$. Simulations cover more than two decades in the diffusion coefficient, corresponding to a Reynolds number (10) which varies from $\text{Re}=600$ to $\text{Re}=158800$, while the large scale dissipation coefficient $\mu$ is fixed. The system is driven by a constant-amplitude forcing with random phases, active within a narrow circular shell in wavevector space centered on $k_{f}$ and with a small width $\Delta k$. This forcing provides constant SPE and VIE injection rates $\varepsilon_{I}$ and $\eta_{I}$ respectively with $\varepsilon_{I}\approx\eta_{I}k_{f}$ since $\Delta k\ll k_{f}$. Specific details on the GPU code performances can be found in (Valadão et al., 2024a).
@@
$A_{1}$ 1024 $642$ $5.5$ $0.0260$- $C_{2}$ 4096 $10600$ $2.9$ $0.0075$ $1130$
$A_{2}$ 1024 $794$ $4.7$ $0.0240$- $C_{3}$ 4096 $15900$ $2.1$ $0.0062$ $1130$
$A_{3}$ 1024 $1060$ $3.8$ $0.0210$- $C_{4}$ 4096 $21200$ $1.7$ $0.0054$ $1130$
$A_{4}$ 1024 $1590$ $2.9$ $0.0180$- $C_{5}$ 4096 $25400$ $1.5$ $0.0049$ $1130$
$B_{1}$ 2048 $2120$ $4.7$ $0.0160$- $D_{1}$ 8192 $31800$ $2.5$ $0.0044$ $9\times 287$
$B_{2}$ 2048 $3180$ $3.5$ $0.0130$ $2270$ $D_{2}$ 8192 $42400$ $2.1$ $0.0039$ $9\times 287$
$B_{3}$ 2048 $6350$ $2.1$ $0.0099$- $D_{3}$ 8192 $63500$ $1.5$ $0.0031$ $9\times 287$
$B_{4}$ 2048 $7940$ $1.8$ $0.0086$ $2270$ $E_{1}$ 16384 $90800$ $2.3$ $0.0026$-
$C_{1}$ 4096 $3970$ $5.9$ $0.0120$- $E_{2}$ 16384 $159000$ $1.6$ $0.0020$-
Table 1: Relevant parameters of the simulation: Reynolds number, diffusive scale $\ell_{\nu}=\nu^{3/4}\varepsilon_{\nu}^{-1/4}$, diffusive time $\tau_{\nu}=\sqrt{\nu/\varepsilon_{\nu}}$, total length of the Lyapunov simulations $T_{\text{tot}}$. Common parameters for all simulations: forcing wavenumber $k_{f}=3.5$ and width $\Delta k_{f}=0.5$, surface potential energy input $\varepsilon_{I}=24$, friction coefficient $\mu=1.0$, characteristic time at the forcing scale $\tau_{f}=\varepsilon_{I}^{-1/3}\ell_{f}^{2/3}=0.51$, maximum resolved wavenumber $k_{\max}=N/3$ (2/3 dealiasing rule).
@@ 3.1 Statistics of the direct cascade
Figure 1 shows the fraction of small-scale dissipation of the inviscid invariants as a function of the Reynolds number of the flow in statistically stationary conditions. Indeed, we find that the small scale relative dissipation of VIE vanishes in the limit of large Re following the prediction (11). On the contrary, since we do not resolve the inverse cascade and $\ell_{\mu}\sim\ell_{f}$, there remains a constant large-scale SPE dissipation $\varepsilon_{\mu}$ for large Re. Therefore, the direct cascade transfers only a fraction of the total injected energy equivalent to $\varepsilon_{\nu}\approx 0.45\varepsilon_{I}$ for $\text{Re}\gtrsim 10^{4}$.
@@
Stationary fluxes of SPE in Fourier space are presented in Fig. 2 for different Reynolds numbers. At moderate $\text{Re}\lesssim 3000$ the fluxes for $k>k_{f}$ decay quickly as a consequence of the viscous dissipation. For large Re, however, a plateau of constant flux emerges at a level corresponding to the viscous dissipation rate $\varepsilon_{\nu}$. We emphasize that SQG turbulence exhibits large fluctuations in the flux of the direct cascade, as studied in details in (Valadão et al., 2024b). These fluctuations arise from the interplay between the accumulation of energy in large-scale structures and intense dissipative events triggered by the formation of filamentary shocks that transfer energy from large to small scales over short time intervals. Thus, very long integrations are necessary to observe the convergence to the constant flux plateau of Fig. 2.
@@
The time-averaged spectra $E(k)$ of SPE are shown in Fig. 3 for all the runs in Table 1. All spectra exhibit power-law behavior, $E(k)\propto k^{-\beta}$ in an intermediate range of scales, which becomes wider as Re increases. We observe that at moderate Re, when a scaling range is already clearly observable, the scaling exponent $\beta$ deviates significantly from the dimensional prediction $5/3$, a feature already reported by previous investigations at comparable Reynolds numbers (Pierrehumbert et al., 1994; Ohkitani & Yamada, 1997; Sukhatme & Pierrehumbert, 2002). Nonetheless, we find that when Re is sufficiently large, $\text{Re}\gtrsim 10^{4}$, the scaling of the dimensional prediction (8) is closely recovered.
@@
In Fig. 4, we plot the dependence of $\xi$ on Re, together with the spectra of Fig. 3 compensated with the expression (12). It is evident that, while for $\text{Re}\lesssim 10^{4}$, the exponent correction $\xi$ depends on Re (approximately as $\text{Re}^{-3/4}$), for larger values of Re, the correction decreases much faster and becomes smaller than $5\%$ for $\text{Re}>2\times 10^{4}$. In this limit, we also observe the convergence of the dimensionless constant to $C_{K}=5.05\pm 0.11$ (see Fig. 3).
@@ 3.2 Predictability of the direct cascade
For the study of chaos and predictability of the direct cascade of SPE, we are interested in computing how two solutions $\theta(\bm{x},t)$ and $\theta^{\prime}(\bm{x},t)$ separate in time on average. In this paper, we consider infinitesimally close solutions so that the average separation rate is given by the maximal Lyapunov exponent of the flow.
Starting from a solution $\theta(\bm{x},t)$ of (1) in a statistically stationary state, we generate a perturbed solution as $\theta^{\prime}(\bm{x},t)=\theta(\bm{x},t)+2\sqrt{\Delta}W({\bm{x}})$ where $W({\bm{x}})$ is a Gaussian random noise with zero mean and unit variance while $\Delta$ is a small parameter. The SPE error $E_{\Delta}$ is defined, for any time, as
$E_{\Delta}(t)={1\over 2}\langle\delta\theta({\bm{x}},t)^{2}\rangle$ (13)
@@
$\gamma_{\tau}(t)={1\over 2\tau}\ln\left({E_{\Delta}(t+\tau)\over\Delta}\right)$ (14)
@@
By repeating the steps (14) and (15) over many time intervals of the same length $\tau$, we obtain a distribution of FTLE along the trajectory. The rescaling procedure (15) ensures the permanence of the perturbation in the exponential growth regime when $\Delta$ and $\tau$ are sufficiently small.
@@
$\gamma_{T}(t)={1\over n}\sum_{k=1}^{n}\gamma_{\tau}(t+k\tau)$ (16)
@@
In general, the distribution of FTLE around the Lyapunov exponent, for sufficiently large $T$, follows the large deviation principle (Vulpiani et al., 2009) which states that
$\rho(\gamma_{T})={1\over N_{T}}e^{-TC(\gamma_{T})}$ (18)
where $N_{T}$ is a normalizing factor and $C(\gamma_{T})$ is the Cramér function, independent on $T$ which, in general, vanishes at $\gamma_{T}=\lambda$ and is positive for $\gamma_{T}\neq\lambda$. For not too large fluctuations, the Cramér function can be approximated by a quadratic form
@@
$\Omega=\lim_{T\to\infty}T\langle(\gamma_{T}-\lambda)^{2}\rangle$ (20)
and it is expected to be $T$-independent in the limit of large $T$.
Figure 5: Fields $\theta(\bm{x})$ and $\delta\theta(\bm{x})$ of Run $D_{3}$ are shown in the left and right panels, respectively. The color code of $\theta(\bm{x})$ shows white/black as the most negative/positive value. The right panel’s color code is based on a log scale $(\propto\log_{10}(|\delta\theta(\bm{x})|))$ to facilitate visualization.
We computed the FTLE for simulations at different Reynolds numbers corresponding to runs $B_{2}$, $B_{4}$, $C_{2}$, $C_{3}$, $C_{4}$, $C_{5}$, $D_{1}$, $D_{2}$, and $D_{3}$ of Table 1. For all runs, we excluded the initial FTLEs in Eq. (14) from the statistics, allowing the perturbation to align with the most unstable direction of the system. For the three cases at the highest Reynolds numbers, we compensated for the increased computational cost by averaging over nine independent, shorter simulations run in parallel, each of which with different realizations of the forcing $f(\bm{x})$ and initial perturbation noise $W(\bm{x})$.
Figure 5 shows a representative realization of the field $\theta(\bm{x})$ along with its corresponding perturbation field $\delta\theta(\bm{x})$. The error seems to accumulate predominantly in filamentary zones between coherent structures. These regions are dominated by small-scale structures formed by energy transfer from larger scales (Pierrehumbert et al., 1994). Since such structures appear intermittently in time, the convergence of the FTLE statistics requires very long simulations, as discussed below.
@@
From Fig. 6, it is evident that the Lyapunov exponent increases with the Reynolds number. By adapting the dimensional arguments developed for 3D NS turbulence (Ruelle, 1979), we expect that $\lambda$ is proportional to the inverse of the smallest dynamical time, i.e., the Kolmogorov time $\tau_{\nu}\equiv(\nu/\varepsilon_{\nu})^{1/2}$
$\lambda\simeq{1\over\tau_{\nu}}\simeq{1\over\tau_{f}}\text{Re}^{1/2}\ .$ (21)
@@
In Fig. 7, we plot the Lyapunov exponents of our simulations as functions of Re. We find that $\lambda$ grows with Re faster than what predicted by (21) and the best fit gives $\lambda\tau_{f}\simeq\text{Re}^{0.7}$ or, equivalently, $\lambda\tau_{\nu}\simeq\text{Re}^{0.2}$. In Fig. 7 we also plot the scaled variance $\Omega$ as a function of Re, which displays a scaling law compatible with that of the Lyapunov exponent $\Omega\tau_{f}\simeq\text{Re}^{0.7}$. This indicates that, in the range of Re investigated here, the ratio $\Omega/\lambda$ is approximately constant ($0.23\pm 0.02$) and that the central part of the Cramér function has a self-similar evolution with Re.
A qualitatively similar behavior has been observed for the Lyapunov exponent of 3D turbulence (Boffetta & Musacchio, 2017; Berera & Ho, 2018; Ge et al., 2023) but with a correction to the dimensional scaling (21) smaller than in the present case. We remark that the origin of such a correction, in both cases, cannot be attributed to intermittent fluctuations in the cascade, which are present also in the SQG cascade of SPE (Valade et al., 2025), since intermittency correction would require an exponent which is smaller than the dimensional one predicted by (21) (Aurell et al., 1996).
Figure 8: Cramér function for the runs at different Re normalized with their corresponding Lyapunov exponent $\lambda$. The dashed line represents the quadratic form $(x-1)^{2}/(2\Omega/\lambda)$ with $\Omega/\lambda=0.23$.
In figure 8, we show the Cramér function $C(\gamma_{T})$ for the different runs obtained from Eq. (18) at sufficiently large values of $T\lambda$. Since $\lambda$ and $\Omega$ scale with Re with the same exponent, we rescale both $\gamma_{T}$ and $C(\gamma_{T})$ by $\lambda$ which, according to the quadratic approximation (19), predicts the collapse to the function $(x-1)^{2}/(2\Omega/\lambda)$. From Fig. 8, we see a remarkable collapse of the Cramér functions at different Re even if the fluctuation for large positive $\gamma_{T}$ is far from the Gaussian prediction (18).

### v2

@@ 1 Introduction
In this work, we consider the surface quasi-geostrophic (SQG) equation, a model that describes the flow governed by the conservation of buoyancy at the surface of a rotating, stratified fluid (Blumen, 1978; Pierrehumbert et al., 1994). Beyond its geophysical relevance, the SQG model has gained attention in the fluid dynamics community due to its striking similarities to three-dimensional (3D) Navier-Stokes (NS) turbulence while keeping some properties of two-dimensional (2D) flows. In particular, SQG has two inviscid quadratic invariants, similarly to 2D turbulence (Celani et al., 2004; Lapeyre, 2017; Valade et al., 2024) with one of the two, the surface potential energy, which displays, in the presence of forcing and dissipation, a direct cascade à la Kolmogorov towards the small scales similar to 3D turbulence (Valadão et al., 2024). Despite this similarity, previous numerical studies (Pierrehumbert et al., 1994; Ohkitani & Yamada, 1997; Sukhatme & Pierrehumbert, 2002) reported that the scaling exponent of the spectrum of the surface potential energy deviates from the Kolmogorov value 5/3-5/3- 5 / 3 predicted by dimensional arguments.
Based on very high-resolution direct numerical simulations of the SQG model, we study the predictability of the direct cascade in relation to the scaling properties of the surface buoyancy field. At large Reynolds numbers, we observe the recovery of the Kolmogorov scaling in the surface potential energy spectrum. Further, we measure the finite-time distribution of the Lyapunov exponent as a function of Re and we find an anomalous scaling law in which the Lyapunov exponent grows faster than what is predicted on dimensional grounds, similar to what was observed in 3D NS turbulence (Boffetta & Musacchio, 2017; Mohan et al., 2017; Berera & Ho, 2018). Despite this anomaly, we find that the distribution of the finite-time Lyapunov exponents follow an almost universal function, independent of the Reynolds number of the flow.
@@ 2 The Surface Quasi-geostrophic model
The SQG model describes the large-scale dynamics of a rapidly rotating, stably stratified flow using a two-dimensional equation for the surface buoyancy field $\theta({\bm{x}},t)$ (Pierrehumbert et al., 1994; Juckes, 1994; Lapeyre & Klein, 2006; Vallis, 2017; Lapeyre, 2017; Siegelman et al., 2022):
$\partial_{t}\theta+{\bf v}\cdot{\bm{\nabla}}\theta=\nu\nabla^{2}\theta-\mu\nabla^{-2}\theta+f\;.$ (1)
The incompressibility condition $\bm{\nabla}\cdot{\bf v}=0$ is enforced through the stream function $\psi(\bm{x},t)$ which defines the velocity field as ${\bf v}({\bm{x}},t)=(-\nabla_{y}\psi,\nabla_{x}\psi)$. The relation between the buoyancy field and the stream function is given by $\psi=\nabla^{-1}\theta$ or, in Fourier space, $\hat{\psi}=\hat{\theta}/k$ (where $k\equiv|\bm{k}|$) and consequently the velocity field can be written in terms of buoyancy field as
@@
from which one observes that $\theta$ has the dimension of a velocity.
@@
The dissipative and forcing terms in the SQG equation represent the effects of scales not resolved by the model, their specific form is somewhat arbitrary (Smith et al., 2002; Lapeyre & Klein, 2006; Burgess et al., 2015). In (1) we introduce a diffusivity $\nu$ and a large-scale damping $\mu$ chosen to be active at, respectively, small and large scales only, while the forcing is active on a narrow range of scales around $\ell_{f}$.
Under these conditions, one expects that a turbulent flow develops with a double cascade phenomenology (Blumen, 1978; Pierrehumbert et al., 1994). Within this scenario, SPE is primarily transferred from the forcing scale to smaller ones ($\ell<\ell_{f}$), producing the direct cascade which is eventually dissipated by viscosity at the diffusive scale $\ell_{\nu}$. Meanwhile, VIE undergoes an inverse cascade, transferring energy to scales larger than the forcing ($\ell>\ell_{f}$) until it is dissipated at the friction scale $\ell_{\mu}$. In the statistically stationary state, the SPE and VIE balances are given by
@@
Under the assumptions of statistical homogeneity and isotropy, Blumen (1978) first predicted the power-law behavior of the energy spectrum $E(k)\equiv\langle|\hat{\theta}({\bm{k}})|^{2}\rangle/2$ for sufficiently large scale separations $\ell_{\nu}\ll\ell_{f}\ll\ell_{\mu}$. In this case, the spectral energy densities follow
@@
Note that in (10) Re is based on $\varepsilon_{I}$ and it is therefore defined a priori. Alternatively, we could use the VIE dissipation $\varepsilon_{\nu}$ resulting in a slightly smaller value of Re, as discussed below.
The scaling laws in (7-8) can also be obtained from the analogous for SQG of the exact four-fifths law of turbulence (Valadão et al., 2024; Valade et al., 2025) and by assuming self-similarity of the statistics (Frisch, 1995). Following this approach, one expects velocity and surface buoyancy to have the same scaling exponent 1/31/31 / 3, which corresponds to that of the velocity field in 3D NS. This observation allows us to adapt to SQG the dimensional arguments developed by Ruelle for the predictability of 3D NS turbulence (Ruelle, 1979). According to the latter, the Lyapunov exponent $\lambda$ characterizing the exponential growth of an infinitesimal perturbation of a solution of (1) should be proportional to the inverse of the smallest dynamical time, i.e., the Kolmogorov time $\tau_{\nu}\equiv(\nu/\varepsilon_{\nu})^{1/2}$
$\lambda\simeq\frac{1}{\tau_{\nu}}\simeq\frac{1}{\tau_{f}}\text{Re}^{1/2}\ .$ (11)
In the following Section, we numerically investigate the scaling properties of the surface buoyancy field and the prediction (11).
@@ 3 Numerical simulations and results
[…]
@@
$A_{1}$ 1024 642642642 5.55.55.5 0.02600.02600.0260 -- $C_{2}$ 4096 106001060010600 2.92.92.9 0.00750.00750.0075 152015201520
$A_{2}$ 1024 794794794 4.74.74.7 0.02400.02400.0240 -- $C_{3}$ 4096 159001590015900 2.12.12.1 0.00620.00620.0062 152015201520
$A_{3}$ 1024 106010601060 3.83.83.8 0.02100.02100.0210 -- $C_{4}$ 4096 212002120021200 1.71.71.7 0.00540.00540.0054 152015201520
$A_{4}$ 1024 159015901590 2.92.92.9 0.01800.01800.0180 -- $C_{5}$ 4096 254002540025400 1.51.51.5 0.00490.00490.0049 152015201520
$B_{1}$ 2048 212021202120 4.74.74.7 0.01600.01600.0160 -- $D_{1}$ 8192 318003180031800 2.52.52.5 0.00440.00440.0044 9×2879$\times$ 2879 × 287
$B_{2}$ 2048 318031803180 3.53.53.5 0.01300.01300.0130 227022702270 $D_{2}$ 8192 424004240042400 2.12.12.1 0.00390.00390.0039 9×2879$\times$ 2879 × 287
$B_{3}$ 2048 635063506350 2.12.12.1 0.00990.00990.0099 -- $D_{3}$ 8192 635006350063500 1.51.51.5 0.00310.00310.0031 9×2879$\times$ 2879 × 287
$B_{4}$ 2048 794079407940 1.81.81.8 0.00860.00860.0086 227022702270 $E_{1}$ 16384 908009080090800 2.32.32.3 0.00260.00260.0026 --
$C_{1}$ 4096 397039703970 5.95.95.9 0.01200.01200.0120 -- $E_{2}$ 16384 159000159000159000 1.61.61.6 0.00200.00200.0020 --
Table 1: Relevant parameters of the simulation: Reynolds number, diffusive scale $\ell_{\nu}=\nu^{3/4}\varepsilon_{\nu}^{-1/4}$, diffusive time $\tau_{\nu}=\sqrt{\nu/\varepsilon_{\nu}}$, total length of the Lyapunov simulations $T_{\text{tot}}$ (999 independent realizations of length $287\tau_{f}$ for the runs $D$). Common parameters for all simulations: forcing wavenumber $k_{f}=3.5$ and width $\Delta k_{f}=0.5$, surface potential energy input $\varepsilon_{I}=24$, friction coefficient $\mu=1.0$, characteristic time at the forcing scale $\tau_{f}=\varepsilon_{I}^{-1/3}\ell_{f}^{2/3}=0.51$, maximum resolved wavenumber $k_{\max}=N/3$ (2/3 dealiasing rule).
@@ 3.1 Statistics of the direct cascade
Figure 1 shows the fraction of small-scale dissipation of the inviscid invariants as a function of the Reynolds number of the flow in statistically stationary conditions. Indeed, we find that the small-scale relative dissipation of VIE vanishes in the limit of large Re following the prediction (12). On the contrary, since we do not resolve the inverse cascade and $\ell_{\mu}\sim\ell_{f}$, there remains a constant large-scale SPE dissipation $\varepsilon_{\mu}$ for large Re. Therefore, the direct cascade transfers only a fraction of the total injected energy equivalent to $\varepsilon_{\nu}\approx 0.45\varepsilon_{I}$ for $\text{Re}\gtrsim 10^{4}$.
@@
[…]
@@
The time-averaged spectra $E(k)$ of SPE are shown in Fig. 3 for all the runs in Table 1. All spectra exhibit power-law behavior, $E(k)\propto k^{-\beta}$ in an intermediate range of scales, which becomes wider as Re increases. We observe that at moderate Re, when a scaling range is already clearly observable, the scaling exponent $\beta$ deviates significantly from the dimensional prediction 5/35/35 / 3, a feature already reported by previous investigations at comparable Reynolds numbers (Pierrehumbert et al., 1994; Ohkitani & Yamada, 1997; Sukhatme & Pierrehumbert, 2002). Nonetheless, we find that when Re is sufficiently large, $\text{Re}\gtrsim 10^{4}$, the scaling of the dimensional prediction (8) is closely recovered.
@@
[…]
@@ 3.2 Predictability of the direct cascade
In this Section, we investigate the predictability of the cascade by computing how two solutions $\theta(\bm{x},t)$ and $\theta^{\prime}(\bm{x},t)$ separate in time on average. We will consider infinitesimally close solutions so that the average separation rate is given by the maximal Lyapunov exponent of the flow.
Starting from a solution $\theta(\bm{x},t)$ of (1) in a statistically stationary state, we generate a perturbed solution as $\theta^{\prime}(\bm{x},t)=\theta(\bm{x},t)+2\sqrt{\Delta}W({\bm{x}})$ where $W({\bm{x}})$ is a Gaussian random white noise with zero mean and unit variance while $\Delta$ is a small parameter. The SPE error $E_{\Delta}$ is defined, for any time, as
$E_{\Delta}(t)=\frac{1}{2}\langle\delta\theta({\bm{x}},t)^{2}\rangle$ (14)
@@
$\gamma_{\tau}(t)=\frac{1}{2\tau}\ln\left(\frac{E_{\Delta}(t+\tau)}{\Delta}\right)$ (15)
@@
By repeating the steps (15) and (16) over many time intervals of the same length $\tau$, we obtain a distribution of FTLE along the trajectory. The rescaling procedure (16) ensures the permanence of the perturbation in the exponential growth regime when $\Delta$ and $\tau$ are sufficiently small (Vulpiani et al., 2009)
@@
$\gamma_{T}(t)=\frac{1}{n}\sum_{k=1}^{n}\gamma_{\tau}(t+k\tau)$ (17)
@@
In general, the distribution of FTLE around the Lyapunov exponent, for sufficiently large $T$, follows the large deviation principle (Vulpiani et al., 2014) which states that
$\rho(\gamma_{T})=\frac{1}{N_{T}}e^{-TC(\gamma_{T})}$ (19)
where $N_{T}$ is a normalizing factor and $C(\gamma_{T})$ is the Cramér function, independent of $T$ which, in general, vanishes at $\gamma_{T}=\lambda$ and is positive for $\gamma_{T}\neq\lambda$ (Boffetta et al., 2002). For not too large fluctuations, the Cramér function can be approximated by a quadratic form
@@
$\Omega=\lim_{T\to\infty}T\langle(\gamma_{T}-\lambda)^{2}\rangle_{\gamma}$ (21)
where the brackets represents average is over the distribution (19) while $\Omega$ is expected to be $T$-independent in the limit of large $T$.
Figure 5: Fields $\theta(\bm{x})$ and $\delta\theta(\bm{x})$ of run $D_{3}$ are shown in the left and right panels, respectively. The color code of the left panel is based on $\theta(\bm{x})/\sup_{\bm{x}}|\theta(\bm{x})|$. The right panel’s color code is based on a log scale $(=\log_{10}(|\delta\theta(\bm{x})|/\sup_{\bm{x}}|\delta\theta(\bm{x})|))$ to facilitate visualization.
We computed the FTLE for simulations at different Reynolds numbers corresponding to runs $B_{2}$, $B_{4}$, $C_{2}$, $C_{3}$, $C_{4}$, $C_{5}$, $D_{1}$, $D_{2}$, and $D_{3}$ of Table 1. For all runs, we excluded from the statistics the initial transient during which the perturbation aligns with the most unstable direction of the system. For the three cases at the highest Reynolds numbers, we compensated for the increased computational cost by averaging over nine independent, shorter simulations run in parallel, each of which with different realizations of the forcing $f(\bm{x})$ and initial perturbation noise $W(\bm{x})$. All rescaling times $\tau$ were kept around the Kolmogorov time scale of the simulation $\tau_{\nu}$, more precisely, between $\tau/\tau_{\nu}\in[0.5,0.8]$ depending on the run.
Figure 5 shows a representative realization of the field $\theta(\bm{x})$ along with its corresponding perturbation field $\delta\theta(\bm{x})$. The errors accumulate predominantly in filamentary zones between coherent structures. These regions are dominated by small-scale structures formed by energy transfer from larger scales (Pierrehumbert et al., 1994). Since such structures appear intermittently in time, the convergence of the FTLE statistics requires very long simulations, as discussed below.
Figure 6: Energy spectra $E(k)$ (full lines) and error spectra $E_{\Delta}(k)$ (dotted line) for the different runs as functions of $k\ell_{\nu}$. The error spectra are computed from the difference field $\delta\theta$ before the rescaling (16) to the initial error.
The distribution of the error field among scales, right before the rescaling, is shown in Fig. 6 where we plot, together with the energy spectra $E(k)$, the error spectra defined in a similar way as $E_{\Delta}(k)=\langle|\delta\hat{\theta}({\bf k})|^{2}\rangle/2$. We observe that for all simulations, the error is concentrated at small scales $k\ell_{\nu}\simeq 0.1$ close to the dissipative range. We remark that the relative magnitude $E_{\Delta}(k)/E(k)$ is very small, ensuring that the perturbation remains in the linear regime.
@@
From Fig. 7, it is evident that the Lyapunov exponent increases with the Reynolds number. As discussed in Section 2, we expect that $\lambda$ follows the Ruelle scaling (11). The latter relies on the spectrum having a Kolmogorov-like power-law behavior. As shown in Fig. 4, the scaling exponent 5/35/35 / 3 is recovered only for $\text{Re}>2\times 10^{4}$. Ruelle’s prediction is therefore expected to hold only for large Re.
@@
In Fig. 8, we plot the Lyapunov exponents of our simulations as functions of Re. We find that $\lambda$ grows with Re faster than what predicted by (11) and the best fit gives $\lambda\tau_{f}\simeq\text{Re}^{0.7}$ or, equivalently, $\lambda\tau_{\nu}\simeq\text{Re}^{0.2}$. Remarkably, the scaling persists also at $\text{Re}<10^{4}$, where the spectrum displays a significant correction to the Kolmogorov exponent 5/35/35 / 3 (see Fig. 4).
In Fig. 8 we also plot the scaled variance $\Omega$ as a function of Re, which displays a scaling law compatible with that of the Lyapunov exponent $\Omega\tau_{f}\simeq\text{Re}^{0.7}$. This indicates that, in the range of Re investigated here, the ratio $\Omega/\lambda$ is approximately constant (0.24±0.020.24$\pm$ 0.020.24 ± 0.02) and that the central part of the Cramér function has a self-similar evolution with Re.
A qualitatively similar behavior has been observed for the Lyapunov exponent of 3D turbulence (Boffetta & Musacchio, 2017; Mohan et al., 2017; Berera & Ho, 2018; Ge et al., 2023) with a correction to the dimensional scaling (11) slightly smaller than in the present case, $\lambda\propto Re^{0.64}$. We remark that the origin of this correction in 3D turbulence is still unclear since it cannot be simply attributed to intermittency. Indeed, the multifractal extension of the dimensional prediction (11) predicts, for 3D turbulence, an exponent smaller than 1/21/21 / 2 (Aurell et al., 1996).
Figure 9: Left: Cramér function for run $C_{4}$ and different times. Right: Cramér functions for different Re at fixed $T\lambda=7$, dashed line represents the quadratic form $(x-1)^{2}/(2\Omega/\lambda)$ with $\Omega/\lambda=0.24$.
[…]
Since $\lambda$ and $\Omega$ scale with Re with the same exponent, we rescale both $\gamma_{T}$ and $C(\gamma_{T})$ by $\lambda$ which, according to the quadratic approximation (20), predicts the collapse to the function $(x-1)^{2}/(2\Omega/\lambda)$.
Figure 9 shows that the quadratic approximation (20) with $\Omega/\lambda=0.24$ describes well the core of $C(\gamma_{T})$ (for $\gamma\simeq\lambda$), while we observe significant deviations for $\gamma_{T}>\lambda$. This means that the situations in which the dynamics of the system is strongly unpredictable (i.e. with $\gamma_{T}\gg\lambda$) are more frequent than what is expected by Gaussian statistics. Nonetheless, it is remarkable that the probability of these extreme deviations seem to be independent of the Reynolds numbers, as shown by the collapse of the right tail of $C(\gamma_{T})$ for different Re.